# Surprisingly fragile: Assessing and addressing prompt instability in multimodal foundation models

### v1

@@ Results
We perform a systematic error analysis to determine possible reasons for the performance improvement on the perturbed prompts (Table 5), comparing the perturbation-trained models to the prompt-trained models. We first embed all the modality data using the ImageBind model (Prompt augmentation), reduce to $D=3$ dimensions with PCA, and then cluster data from each modality via cosine similarity of the embeddings. The clustering was implemented with a hierarchical DBSCAN with default hyperparameters (minimum cluster size 5, automatic algorithm selection) which identified $K={15,11,12}$ clusters for audio, image, and video modalities respectively. We identify the dominant concept in each cluster based on the associated prompts, and identify clusters with an especially high improvement from the perturbation-trained models as compared to the prompt-trained models. The high-improvement clusters are shown in Table 7, e.g., the image cluster related to street views shows a roughly 250% improvement from prompt-trained to perturbation-trained (0.3713 to 0.9050).
@@
For audio, the example cluster represents spoken explanations of religious organizations, such as the UMC (United Methodist Church). This topic of data likely requires more complicated and contextualized answers to questions, which the perturbation-trained models seem able to provide (e.g. OFASys response, “largest protestant”) while the models trained only on original prompts produce more short and ambiguous responses (OFASys, “national”). For image, the example cluster contains scenes of home activities, including a typical kitchen scene. These data include prompts that ask about basic object properties such as quantity and material type, and the perturbation-trained models appear more able to understand the intent of the prompts as they relate to properties such as “substance” (e.g., Unified-IO, “marble”). For video, the example cluster includes clips of people facing the camera, mainly performing tasks with tools such as applying makeup and combing their hair. The prompts for this data ask for further descriptions of the actions shown, and the prompt-trained models have difficulty understanding the full context of differently-phrased actions (e.g., OFASys, “dry”) as compared to the perturbation-trained models are able to provide a full description of the actions (OFASys, “The woman in the video combed her hair with a brush.”).
We show the top-3 clusters per modality in Table 8 to show more example prompts and aggregate model scores. The improvement across highly diverse topics and contexts shows that the perturbation-trained models’ improvements are not spurious (e.g., only applying to a specific type of QA task) but reflect an improvement in general reasoning ability.
Conclusion
@@
This study showed that multimodal foundation models (MFMs) show a surprising degree of instability in response to prompt perturbations, and that the instability can be mitigated through fine-tuning on additional prompt perturbations, regardless of whether the perturbations are close to the original prompt or modality data. The finding about re-training on prompt perturbations is surprising but suggests that MFM developers should focus more on diversity of prompts rather than prompt similarity to the original data, which appears to be less relevant for models during training. Lastly, the error analysis shows systematic improvements in specific clusters of data in different modalities, e.g. in videos that depict personal grooming, and this suggests that the additional training helps reinforce modality-specific properties to the model that leads to greater robustness.
@@
Future work may investigate more complicated paradigms for increasing model robustness, including direct preference optimization to determine the relative value of the different perturbations for model training (rafailov2024; zhao2024). In this work we only considered perturbations of the text data, but future work should test model robustness in response to perturbations of the modality data, e.g., adding additional noise to images (lee2020). Lastly, future work should consider more structured/explainable methods for determining similarity between modality data and text: instead of ImageBind, one may use the texts output from a well-trained image caption model to compare against a prompt perturbation to determine if the perturbation has “drifted” too far from the image content.
Acknowledgments
This project was funded by NA-22…TODO: funding details
Limitations
The prompt perturbations in this work are meant to emulate real-world variation in prompts due to differences in user input. However, this is a strong assumption that may not hold in practice, and we cannot guarantee that the perturbations generated in this study are a perfect match for user input. We hope that future work can investigate this mismatch more closely and determine the best proxy for variation in real-world prompts. Ideally, this work would provide a lower bound for the possible diversity in prompt space due to differences among system users, considering that human behavior with respect to prompting LLMs can be somewhat unpredictable (kim2024).
Appendix A Results by prompt perturbation type
We further analyze the aggregate prompt perturbation results (Table 5) by sub-dividing them by prompt perturbation type, i.e. joint similarity, modality-similarity, and text-similarity. The results per-type are shown in separate tables: text-similarity prompt results in Table 9, modality-similarity prompt results in Table 10, and joint similarity prompt results Table 11.
The prompt-perturbation models have higher accuracy for their respective perturbation type, e.g. for the text-similarity prompts the model with highest performance was the model trained on text-similarity prompts. This suggests that the prompt perturbation filtering processes tend to generate self-similar prompts, which leads to more consistent performance from the models that are trained on those filtered prompts when applied to those self-similar prompts. The self-similarity is apparent even in the example prompts in LABEL:fig:prompt_sample_method, where the text-similarity prompt perturbations focus on the act of “holding” rather than the object being held or the person doing the holding.
OFASys Unified-IO
Modality Model BLEU ROUGE BERT BLEU ROUGE BERT
Audio No training 0.0450 0.1337 0.8169 0.0145 0.0469 0.8170
Original prompts 0.1279 0.1630 0.8774 0.0449 0.0709 0.8521
Joint similarity 0.4486 0.4797 0.9225 0.2202 0.2399 0.8842
Modality-similarity 0.4518 0.4853 0.9233 0.2350 0.2519 0.8859
Text-similarity 0.4941 0.5239 0.9274 0.3158 0.3320 0.8976
Random sample 0.4138 0.4492 0.9186 0.2432 0.2604 0.8850
Image No training 0.3947 0.4045 0.9644 0.3827 0.3920 0.9308
Original prompts 0.5026 0.5107 0.9699 0.4083 0.4193 0.9423
Joint similarity 0.7308 0.7339 0.9807 0.6673 0.6731 0.9824
Modality-similarity 0.7253 0.7287 0.9802 0.6848 0.6891 0.9837
Text-similarity 0.7767 0.7809 0.9840 0.7383 0.7432 0.9856
Random sample 0.7232 0.7280 0.9806 0.6870 0.6921 0.9835
Video No training 0.0045 0.0314 0.8170 0.0062 0.0364 0.8177
Original prompts 0.0230 0.0745 0.8250 0.1588 0.2967 0.8562
Joint similarity 0.4182 0.6011 0.9451 0.4675 0.6224 0.9463
Modality-similarity 0.4177 0.6037 0.9452 0.4892 0.6326 0.9478
Text-similarity 0.4402 0.6204 0.9478 0.6097 0.7179 0.9599
Random sample 0.4167 0.6009 0.9455 0.4681 0.6159 0.9461
Table 9: Model accuracy on text-similarity perturbations.
OFASys Unified-IO
Modality Model BLEU ROUGE BERT BLEU ROUGE BERT
Audio No training 0.0446 0.1335 0.8167 0.0126 0.0456 0.8175
Original prompts 0.1376 0.1787 0.8750 0.0464 0.0768 0.8529
Joint similarity 0.4300 0.4646 0.9190 0.2178 0.2404 0.8830
Modality-similarity 0.4670 0.4984 0.9241 0.2825 0.2978 0.8923
Text-similarity 0.4523 0.4867 0.9213 0.2366 0.2571 0.8854
Random sample 0.4123 0.4499 0.9175 0.2201 0.2402 0.8811
Image No training 0.3753 0.3854 0.9642 0.3439 0.3538 0.9266
Original prompts 0.4602 0.4696 0.9653 0.3770 0.3868 0.9442
Joint similarity 0.6971 0.7013 0.9795 0.6330 0.6374 0.9786
Modality-similarity 0.7394 0.7431 0.9826 0.7195 0.7222 0.9849
Text-similarity 0.6913 0.6980 0.9805 0.6322 0.6374 0.9772
Random sample 0.6902 0.6959 0.9806 0.6391 0.6439 0.9802
Video No training 0.0041 0.0292 0.8168 0.0068 0.0360 0.8172
Original prompts 0.0219 0.0694 0.8246 0.1516 0.2843 0.8610
Joint similarity 0.3980 0.5859 0.9429 0.4501 0.6041 0.9439
Modality-similarity 0.4132 0.6007 0.9449 0.5710 0.6863 0.9558
Text-similarity 0.3995 0.5886 0.9433 0.4706 0.6188 0.9462
Random sample 0.3958 0.5858 0.9430 0.4403 0.5945 0.9430
Table 10: Model accuracy on modality-similarity perturbations.
OFASys Unified-IO
Modality Model BLEU ROUGE BERT BLEU ROUGE BERT
Audio No training 0.0448 0.1336 0.8168 0.0154 0.0496 0.8169
Original prompts 0.1232 0.1585 0.8755 0.0439 0.0725 0.8533
Joint similarity 0.4625 0.4949 0.9247 0.2597 0.2739 0.8883
Modality-similarity 0.4379 0.4678 0.9210 0.2268 0.2438 0.8824
Text-similarity 0.4474 0.4802 0.9211 0.2353 0.2543 0.8843
Random sample 0.3959 0.4292 0.9154 0.2292 0.2461 0.8823
Image No training 0.3936 0.4047 0.9636 0.3497 0.3570 0.9289
Original prompts 0.4556 0.4630 0.9670 0.3764 0.3868 0.9362
Joint similarity 0.7399 0.7447 0.9805 0.7223 0.7263 0.9837
Modality-similarity 0.7027 0.7075 0.9798 0.6711 0.6757 0.9795
Text-similarity 0.6871 0.6937 0.9782 0.6396 0.6455 0.9780
Random sample 0.6908 0.6961 0.9788 0.6649 0.6684 0.9805
Video No training 0.0041 0.0287 0.8169 0.0067 0.0352 0.8171
Original prompts 0.0225 0.0737 0.8259 0.1496 0.2871 0.8628
Joint similarity 0.4106 0.5958 0.9439 0.5379 0.6669 0.9526
Modality-similarity 0.3941 0.5866 0.9424 0.4536 0.6043 0.9440
Text-similarity 0.3947 0.5836 0.9429 0.4587 0.6099 0.9447
Random sample 0.3932 0.5805 0.9424 0.4499 0.6003 0.9438
Table 11: Model accuracy on joint similarity perturbations.

### v2

@@ 4.3 Error analysis
We perform a systematic error analysis to determine possible reasons for the performance improvement on the prompt perturbations (accuracy shown in Table 5), and we compare the perturbation-trained models to the prompt-trained models. We first embed all the modality data using the same ImageBind model used for prompt perturbation filtering (3.1), reduce to $D=3$ dimensions with PCA to group dimensions with similar semantics, and then cluster data from each modality via cosine similarity of the reduced embeddings. The clustering is implemented with a hierarchical DBSCAN with default hyperparameters (minimum cluster size 5, automatic algorithm selection) which identified K=15,11,12K={15,11,12} clusters for audio, image, and video modalities respectively. We identify the dominant concept in each cluster based on the associated prompts, and identify clusters with an especially high improvement from the perturbation-trained models as compared to the prompt-trained models.
@@
Examples of high-improvement clusters are shown in Table 7. For audio, the example cluster represents spoken explanations of religious organizations, such as the UMC (United Methodist Church). This topic of data likely requires more complicated and contextualized answers to questions, which the perturbation-trained models seem able to provide (e.g. OFASys response, “largest protestant”) while the models trained only on original prompts produce more short and ambiguous responses (OFASys, “national”). For image, the example cluster contains scenes of home activities, including a typical kitchen scene. These data include prompts that ask about basic object properties such as quantity and material type, and the perturbation-trained models appear more able to understand the intent of the prompts as they relate to properties such as “substance” (e.g., Unified-IO, “marble”). For video, the example cluster includes clips of people facing the camera, mainly performing tasks with tools such as applying makeup and combing their hair. The prompts for this data ask for further descriptions of the actions shown, and the prompt-trained models have difficulty understanding the full context of differently-phrased actions (e.g., OFASys, “dry”) as compared to the perturbation-trained models that provide a complete description of the actions (OFASys, “The woman in the video combed her hair with a brush.”).
We show more examples of per-cluster prompts and quantitative differences in model performance in Appendix, e.g., the image cluster related to street views shows a roughly 250% improvement from prompt-trained to perturbation-trained (0.3713 to 0.9050).
5 Conclusion
@@
This study showed that multimodal foundation models (MFMs) show a surprising degree of instability in the image domain in response to prompt perturbations (4.1). However, the instability can be reduced through fine-tuning on additional prompt perturbations, regardless of whether the perturbations are close in semantic space to the original text data or modality data (4.2). The finding about re-training on different sub-sets of prompt perturbations is surprising but suggests that MFM developers should focus more on diversity of prompts rather than optimizing a specific method for prompt sampling, because the choice of method appears less relevant for downstream model performance. Lastly, the error analysis shows systematic improvements in specific clusters of data in different modalities, e.g. in videos that depict personal grooming (4.3). This suggests that training on prompt perturbations helps reinforce modality-specific properties (e.g., fine-grained visual details in grooming behavior), that in turn leads to greater robustness in those modalities.
@@
This study provides just one approach to tackle the larger problem of model robustness, which future work should continue to address with different methods. For one, future studies may investigate more model-driven paradigms for increasing robustness, including direct preference optimization to determine the relative value of the different perturbations for model training rafailov2024; zhao2024. In this work we only considered perturbations of the text data, but MFMs should also be tested with respect to perturbations of the modality data, e.g., adding additional noise to images lee2020. Lastly, future work should consider more structured or explainable methods for determining similarity between modality data and text. E.g., one might use the text output from a well-trained image caption model to compare against a prompt perturbation to determine if the perturbation has “drifted” too far from the image content.
6 Acknowledgments
This work was supported by the NNSA Office of Defense Nuclear Nonproliferation Research and Development, U.S. Department of Energy, and Pacific Northwest National Laboratory, which is operated by Battelle Memorial Institute for the U.S. Department of Energy under Contract DEAC05–76RLO1830. This article has been cleared by PNNL for public release as PNNL-SA-201679.
Appendix A Experiments on modern multimodal models
We conduct additional experiments to validate the perturbation-related results on modern multimodal models (cf. section˜4.1). The following modern models are tested:
  • Phi-4 abouelenin2025phi;
  • Gemma-3n team2025gemma.
We run the same prompt perturbations through the other models and largely replicate the earlier results in which model performance decreases when faced with text perturbations section˜A.1. We then perturb the modality-specific data by adding noise, swapping sub-sections of data, and masking sub-sections of data section˜A.2, and we find that these perturbations do not consistently decrease or increase performance, which suggests that the main problem with multimodal model instability relates to text-related perturbations.
A.1 Prompt perturbations
We show the results of the prompt perturbations in Table 8. As seen before, the models’ performance on prompt perturbations declines significantly regardless of modality, particularly in the video modality where BLEU scores drop by up to 20% for Gemma and up to 28% for Phi-4. Interestingly, for some perturbations in the image modality for Phi-4 we see increased performance. We attribute this difference in results to a qualitative change in Phi-4 behavior as a result of receiving perturbed prompts. Normally Phi-4 output is verbose and contains superfluous information about the data provided, but when providing the model with noisy data it appears that the responses are often more concise, resulting in higher text overlap scores. Why this only impacts images may be an artifact of the training for Phi-4 that prioritized long detailed captions on images as opposed to other modalities where the fine-tuning may have involved more concise output.
Gemma-3n Phi-4
Modality Prompt BLEU ROUGE BERT BLEU ROUGE BERT
Audio Original 0.08460 (0.00520) 0.18980 (0.00938) 0.82617 (0.00185) 0.21653 (0.01333) 0.34213 (0.01566) 0.86562 (0.00302)
LLM-Paraphrase 0.05259 (0.00125) 0.11038 (0.00242) 0.82056 (0.00054) 0.11954 (0.00321) 0.19136 (0.00403) 0.84177 (0.00082)
Back-translation 0.06959 (0.00332) 0.15202 (0.00637) 0.82510 (0.00125) 0.13939 (0.00738) 0.22462 (0.00963) 0.85045 (0.00185)
Paraphrase 0.06521 (0.00280) 0.14177 (0.00534) 0.82206 (0.00107) 0.15964 (0.00765) 0.24218 (0.00866) 0.84984 (0.00173)
Image Original 0.03542 (0.00244) 0.08650 (0.00522) 0.81276 (0.00098) 0.36079 (0.02548) 0.39153 (0.02451) 0.90119 (0.00478)
LLM-Paraphrase 0.02649 (0.00072) 0.06486 (0.00151) 0.81270 (0.00030) 0.34425 (0.00840) 0.35425 (0.00826) 0.91033 (0.00161)
Back-translation 0.03180 (0.00205) 0.07802 (0.00421) 0.81208 (0.00081) 0.40713 (0.02290) 0.42287 (0.02230) 0.91883 (0.00413)
Paraphrase 0.03189 (0.00150) 0.07864 (0.00314) 0.81263 (0.00058) 0.38722 (0.01598) 0.39820 (0.01560) 0.91072 (0.00297)
Video Original 0.37934 (0.00659) 0.45289 (0.00900) 0.89921 (0.00119) 0.56810 (0.01209) 0.46234 (0.01046) 0.92176 (0.00158)
LLM-Paraphrase 0.31094 (0.00189) 0.35088 (0.00250) 0.88856 (0.00036) 0.40677 (0.00375) 0.33885 (0.00308) 0.90074 (0.00062)
Back-translation 0.32404 (0.00549) 0.35879 (0.00695) 0.88985 (0.00100) 0.42460 (0.00960) 0.33502 (0.00764) 0.90196 (0.00138)
Paraphrase 0.34445 (0.00399) 0.39269 (0.00534) 0.89255 (0.00078) 0.46408 (0.00738) 0.38279 (0.00619) 0.90795 (0.00113)
Table 8: Performance of baseline models (no additional training) on original and perturbed prompts; standard error in parentheses.
A.2 Modality perturbations
A natural question from the prompt perturbation experiments would be whether the same effect holds for perturbations on the modality-specific data, e.g. image noising. We experiment with three different types of modality perturbations:
  • Mask: randomly mask 10% of the data, in regularly-shaped 2-D squares (for image and video) or in 1-D lines (for audio);
  • Noise: randomly add Gaussian noise to the data (normalized to have the same mean and variance as original data);
  • Swap: randomly swap 10% of the data, in regularly-shaped 2-D squares (for image and video) or in 1-D lines (for audio).
We run the evaluation on a smaller sample of data due to memory constraints (N=1000 per modality). The results are shown in Table 9, and we see largely similar performance across modalities regardless of perturbation condition. The primary exception is the improved performance on modality perturbations in the image modality for Phi-4, as seen in the prompt perturbation results. In nearly all cases, we can conclude that modality data perturbations pose much less of a problem to multimodal model performance as compared to prompt perturbations.
Gemma-3n Phi-4
Modality Perturbation BLEU ROUGE BERT BLEU ROUGE BERT
Audio Original 0.08568 (0.00480) 0.19375 (0.00903) 0.82850 (0.00178) 0.21653 (0.01333) 0.34213 (0.01566) 0.86562 (0.00302)
Mask 0.08670 (0.00508) 0.19385 (0.00928) 0.82680 (0.00177) 0.20163 (0.01144) 0.32878 (0.01450) 0.86228 (0.00280)
Noise 0.08864 (0.00520) 0.19667 (0.00966) 0.82759 (0.00187) 0.21017 (0.01217) 0.34339 (0.01566) 0.86528 (0.00289)
Swap 0.08132 (0.00501) 0.18277 (0.00966) 0.82553 (0.00190) 0.19848 (0.01079) 0.33010 (0.01483) 0.86020 (0.00273)
Image Original 0.03622 (0.00248) 0.08974 (0.00502) 0.81352 (0.00095) 0.36079 (0.02548) 0.39153 (0.02451) 0.90119 (0.00478)
Mask 0.03446 (0.00236) 0.08462 (0.00506) 0.81289 (0.00094) 0.63760 (0.02723) 0.63353 (0.02710) 0.96650 (0.00378)
Noise 0.03534 (0.00228) 0.08466 (0.00486) 0.81361 (0.00092) 0.60103 (0.02767) 0.59879 (0.02741) 0.95801 (0.00416)
Swap 0.03342 (0.00230) 0.08149 (0.00499) 0.81341 (0.00094) 0.64869 (0.02698) 0.64401 (0.02683) 0.96906 (0.00373)
Video Original 0.37991 (0.00657) 0.44466 (0.00926) 0.89774 (0.00121) 0.56810 (0.01209) 0.46234 (0.01046) 0.92176 (0.00158)
Mask 0.37495 (0.00681) 0.44338 (0.00918) 0.89778 (0.00124) 0.57727 (0.01168) 0.46501 (0.01025) 0.92336 (0.00150)
Noise 0.37061 (0.00643) 0.44293 (0.00907) 0.89700 (0.00118) 0.58019 (0.01263) 0.46943 (0.01065) 0.92325 (0.00162)
Swap 0.38795 (0.00639) 0.46090 (0.00911) 0.89698 (0.00127) 0.57764 (0.01186) 0.46758 (0.01031) 0.92373 (0.00145)
Table 9: Performance of baseline models (no additional training) on original and perturbed modality data; standard error in parentheses.